the cost of cakes with double egg yolks cost 50 cents more than those with only a single egg yolk. Mrs.Jim bought 6 cakes with double egg yolks and 5 with only a single yolk. if she paid $36 for 11 cakes, how much does a cake with double egg yolk cost?
Answer by TimothyLamb(4379)

(B+50)x + By = C
6(B+50) + 5B = 3600
6B + 300 + 5B = 3600
11B = 3300
B = 300
Answer:
double yolk: $3.50
single yolk: $3.00
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Convert fractions, decimals, and percents:
https://sooeet.com/math/fraction-decimal-percent.php
---
Calculate and graph the linear regression of any data set:
https://sooeet.com/math/linear-regression.php